# SQ-GAN: Semantic Image Communications Using Masked Vector Quantization

### v1

@@ IV-D Fine-tuning $G$
The final step involves fine-tuning the entire network $G$ by freezing the parameters of $G_{\mathbf{s}}$ and updating only those of $G_{\mathbf{x}}$. This fine-tuning addresses the scenario where the original SSM $\mathbf{s}$ is unavailable at the receiver, that is when the SQ-GAN is used for Semantic Communication (SC). Notice that the fine-tuning step is essential since the quality of $\hat{\mathbf{s}}$ is influenced by the masking fraction $m_{\mathbf{s}}$, and this dependency is missed in the separate training steps described above. The loss function and the semantic-aware discriminator $D_{disc}^{\mathbf{x}}$ remain identical to those used in the training of $G_{\mathbf{x}}$. Fine-tuning is conducted over 100 epochs with early stopping to prevent overfitting. During each iteration, both masking fractions $m_{\mathbf{x}}$ and $m_{\mathbf{s}}$ are randomly selected from the same distribution. The Adam optimizer is employed with a learning rate of $10^{-4}$ and a batch size of 8.
@@ V Results
This section focuses on the performances of the proposed SQ-GAN evaluated using 500 pairs of images and SSMs from the validation set of the Cityscape dataset [25].
In Section V-A, we illustrate the function and role of the SAMM in selecting features of $\mathbf{x}$ and $\mathbf{s}$ that are semantically relevant. Additionally, we also show the effect of the masking fractions on the final reconstruction. In Section V-B the performance of the proposed SQ-GAN is compared with state of the art compression algorithms like BPG and JPEG2000.
@@ V-A SAMM function
Figure 7: Visual representation of the latent tensor selection of the SAMM projected in the image and SSM domain. In both cases the masking has been fixed to $m_{\mathbf{x}}=m_{\mathbf{s}}=0.20$ and the region considered semantically relevant are shown on the right.
@@
Figure 8: Visual comparison between the same image and SSM at different masking reactions $m_{\mathbf{x}}$ and $m_{\mathbf{s}}$. The original image and SSM are shown on the left. The upper row shows the reconstructed $\hat{\mathbf{x}}$ and the generated SSM using the SOTA SSModel INTERN-2.5 [21]. The bottom row shows the reconstructed SSM $\hat{\mathbf{s}}$. All pairs $(\hat{\mathbf{x}},\hat{\mathbf{s}})$ are obtained at $0.05$BPP.
SQ-GAN adopts various techniques to force the model to correctly reconstruct the relevant regions of the image. The weighted loss function based on the semantic classes, the semantic-aware discriminator $D_{disc}^{\mathbf{x}}$, and the semantic relevant classes enhancement data augmentation have been designed with this idea in mind. Their scope is to guide SAMM to identify and select the relevant latent vectors $\mathbf{z}_{k}^{\mathbf{x}}$ and $\mathbf{z}_{k}^{\mathbf{s}}$.
Since the selection is performed on the latent tensor, to understand the effect of SAMM, we project back from the latent tensor to the image (pixel) domain [19]. Fig. 7 shows such a projection for a specific example.  [^2] The projection showed in figure 7 was implemented by leveraging the convolutional structure of the two encoders that allows to trace back the spatial correlation between the latent tensor and the original frame. In the left column the original $\mathbf{x}$ and $\mathbf{s}$ are represented, while the right column shows which are the regions associated to the latent vectors that SAMM considers more relevant. In this example the masking fractions are chosen as $m_{\mathbf{x}}=m_{\mathbf{s}}=0.20$.
It is immediately evident that $G_{\mathbf{x}}$ and $G_{\mathbf{s}}$ consider different regions as relevant. The SAMM in $G_{\mathbf{s}}$ focuses on regions with the most change in semantic classes. The street, building and sky require very few associated latent vectors to be represented. On the contrary the parts of the SSM containing relevant classes are strongly preferred.
In a similar way, the SAMM in $G_{\mathbf{x}}$ shows a strong preference for relevant classes like cars and people. However, it also focuses on areas previously ignored, such as the street and buildings. In fact, thanks to the conditioning on the SSM, the sub-network $G_{\mathbf{x}}$ knows the location of every object and their shape and can focus on different aspects like colors and textures. For this reason the streets and sky will require some latent vectors to reconstruct the colors correctly. However, most of the latent vectors are selected from the truly relevant regions.
Figure 9: SSM retention evaluated between the true $\mathbf{s}$ and the reconstructed $\hat{\mathbf{s}}$ with the mean Intersection over Union (mIoU) metric as a function of the masking fraction $m_{\mathbf{s}}$. As the masking fraction $m_{\mathbf{s}}$ increases the network $G_{\mathbf{s}}$ is able to better reconstruct the SSM. However, the increase of performances reaches a plateau from $m_{\mathbf{s}}\geq 0.2$ ($BPP_{\mathbf{s}}=0.011$BPP).
Figure 10: LPIPS evaluated between $\mathbf{x}$ and $\hat{\mathbf{x}}$ as the masking fractions $m_{\mathbf{s}}$ e $m_{\mathbf{x}}$ vary.
Original $0.019$BPP $0.038$BPP $0.078$BPP
@@
The column on the left shows the original $\mathbf{x}$ and $\mathbf{s}$. The other columns show on top the reconstructed $\hat{\mathbf{x}}$ and the generated SSM obtained form $\hat{\mathbf{x}}$ via the INTERN-2.5 SSModel [21]. The bottom row shows the reconstructed $\hat{\mathbf{s}}$. These examples are obtained by fixing the compression level to a total amount of $0.05$ BPP and by letting the masking fractions vary, such that $m_{\mathbf{x}}+m_{\mathbf{s}}=1.1$ (constant).
Conversely, the influence of $m_{\mathbf{x}}$ on the overall quality shows expected behavior. Increasing $m_{\mathbf{x}}$ enhances image fidelity, particularly for non-relevant details. Relevant features are prioritized and reconstructed effectively even at low $m_{\mathbf{x}}$, but finer details, such as building windows, are better retained when $m_{\mathbf{x}}=0.95$. This highlights the model’s capability to prioritize relevant features before addressing non-relevant ones.
The impact of $m_{\mathbf{s}}$ on the mIoU is illustrated in Fig. 10. It is useful to notice that, for $m_{\mathbf{s}}\geq 0.55$, the value of mioU remains almost constant denoting that, above that value, additional latent vectors do not add relevant semantic features. For very low values of $m_{\mathbf{s}}$ the performance is poor, but, as soon as $m_{\mathbf{s}}\geq 0.20$, the model is able to reconstruct $\hat{\mathbf{s}}$ with an acceptable level of semantic retention. To give a term of comparison, the value of $m_{\mathbf{s}}=0.20$ corresponds to a compression of $BPP_{\mathbf{s}}=0.011$ BPP. This means that at this low value of BPP the model is already able to preserve valuables details in the SSM. Increasing $m_{\mathbf{s}}$ beyond the value of 0.55 (equivalent to $BPP_{\mathbf{s}}=0.025$ BPP) does not provide further improvement. We interpret this saturation as follow: beyond this value, all semantically relevant latent vectors have been selected and quantized. Adding other vectors will only increase the amount of redundant information, without improving the reconstructed SSM. This behavior is quite different from what observed and discussed previously about $m_{\mathbf{x}}$, where even for values close to $1$, the improvements were still visible on the final output.
Figure 12: Performance comparison between BPG, JPEG2000 and SQ-GAN in term of semantic metrics and the classic pixel-by-pixel PSNR.
It is also interesting to see how both $m_{\mathbf{x}}$ and $m_{\mathbf{s}}$ affect $\hat{\mathbf{x}}$. For this purpose, in Fig. 10 we report the value of LPIPS to measure the distance between $\mathbf{x}$ and $\hat{\mathbf{x}}$. As expected, the performance along the $m_{\mathbf{s}}$ axis has a strong influence on the output only for values of $m_{\mathbf{s}}\leq 0.55$. Instead, the influence of $m_{\mathbf{x}}$ is observed along the whole range from 0 to 1. This is consistent with the visual results shown in Fig. 8.
@@ V-B Visual Results and Comparisons with SOTA image compression
In this section the results of the proposed SQ-GAN are compared with SOTA compression algorithms BPG and JPEG2000. A visual comparison between BPG and SQ-GAN is shown in Fig. 11.
The top row represents the reconstructed image $\hat{\mathbf{x}}$ obtained with the proposed SQ-GAN and the associated generated SSM via the INTERN-2.5 SSModel. The bottom row shows the reconstructed image obtained by using the BPG algorithms and the associated generated SSM.
We notice that SQ-GAN can reach lower values of BPP. In particular, in this case BPG cannot compress images at BPP lower than $0.038$. Furthermore, while BPG uses precious resources to reconstruct semantically irrelevant details such as the windows of the buildings, SQ-GAN focuses on the relevant parts. This is further evidenced by the amount of semantic retention of the generated SSM. For example, the cyclist riding the bike is still visible and correctly identified by SSModel, at all levels of BPP in Fig. 11. In contrast, the cyclist completely disappears in the generated SSM out of the BPG reconstructed image at all levels of BPP in Fig. 11.
@@
where $\mathbf{X}$ and $\hat{\mathbf{X}}$ are the sets of real and reconstructed images, respectively; $\mu_{\phi}(\mathbf{X})$ and $\Sigma_{\phi}(\mathbf{X})$ are the mean and covariance of the features extracted from the real images $\mathbf{X}$ using a pre-trained Inception-v3 model $\phi$ [35]; $\mu_{\phi}(\hat{\mathbf{X}})$ and $\Sigma_{\phi}(\hat{\mathbf{X}})$ are the corresponding statistics from the reconstructed images $\hat{\mathbf{X}}$. Rather than directly comparing images pixel by pixel, the FID compares the distributions of the latent representations of the original and reconstructed images obtained using a convolutional neural network (Inception v3) trained for image recognition. From its definition, it is a parameter that is robust against translation, rotation, or change of scale of relevant objects. A low value of FID indicates that a DNN operating over the two batches of original and reconstructed images extract representations that are statistically equivalent.
It is now useful to compare SQ-GAN with state-of-the-art image encoders, like JPEG2000, and BPG, using the following four metrics: FID, LPIPS, Peak Signal-to-Noise Ratio (PSNR), and mIoU. The first three metrics are evaluated between the original $\mathbf{x}$ and the reconstructed images $\hat{\mathbf{x}}$, while the mIoU value is evaluated between $\mathbf{s}$ and the SSM $\tilde{\mathbf{s}}$ generated via the SOTA INTERN-2.5 SSModel. The comparison is reported in Fig. 12, where we show the above metrics vs. the BPP value.
From Fig. 12, SQ-GAN significantly outperforms JPEG2000 and BPG, in terms of parameters that somehow reflect a sort of semantic distortion, like, LPIS, FID, and mIoU. At the same time, SQ-GAN is outperformed by BPG in terms of PSNR, i.e. a standard metric that makes a pixel-by-pixel comparison. Furthermore, SQ-GAN is able to reach very low BPP, in the order of $0.01$ BPP, with still relatively good values of the semantic metrics FID, LPIPS, and mIoU. It is interesting to notice that, at $0.038$ BPP the BPG algorithm can reconstruct images with a Peak Signal-to-Noise Ratio (PSNR) of 26, but it fails to preserve the SSM. As shown in Fig. 11, while BPG reconstructs the the buildings’ windows more accurately, only the first two cars are detected and the rest of the objects are lost.
In summary, Fig. 12 shows the main feature of the proposed method: In all applications where the task is to compress in order to preserve the semantic content of the image, without necessarily being able to reconstruct images that are very similar to the original ones at the pixel level, SQ-GAN significantly outperforms state-of-the art image encoders.
@@ VI Conclusions
This work presents SQ-GAN, a new approach for simultaneous encoding of images and their semantic maps, aimed at preserving the image semantic content, even at very low values of bit per pixel. The comparison of SQ-GAN with state-of-the-art encodes has been carried out using a set of performance indicators, like FID, LPIPS, and mIoU, that somehow measure the semantic distortion between the original image (or semantic map) and the reconstructed image (or semantic map). Using these performance indicators, extensive evaluations on the Cityscapes dataset show that SQ-GAN significantly outperforms JPEG2000 and BPG, in terms of rate-(semantic) distortion tradeoff. At the same time, SQ-GAN is clearly outperformed by BPG in terms of conventional pixel-based indicators, such as the PSNR. In summary, the method is appealing whenever compression is performed having a clear task in mind, motivating the compression like, for example, object recognition from images collected from remote sensors.
Further investigations are necessary to simplify the coding strategy and to generalize the approach to video coding.

### v2

@@ IV-D Fine-tuning $G$
The final step involves fine-tuning the entire network $G$ by freezing the parameters of $G_{\mathbf{s}}$ and updating only those of $G_{\mathbf{x}}$. This fine-tuning addresses the scenario where the original SSM $\mathbf{s}$ is unavailable at the receiver, that is, when the SQ-GAN is used for Semantic Communication (SC). Notice that the fine-tuning step is essential since the quality of $\hat{\mathbf{s}}$ is influenced by the masking fraction $m_{\mathbf{s}}$, and this dependency is missed in the separate training steps described above. The loss function and the semantic-aware discriminator $D_{disc}^{\mathbf{x}}$ remain identical to those used in the training of $G_{\mathbf{x}}$. Fine-tuning is conducted over 100 epochs with early stopping to prevent overfitting. During each iteration, both masking fractions $m_{\mathbf{x}}$ and $m_{\mathbf{s}}$ are randomly selected from the same distribution. The Adam optimizer is employed with a learning rate of $10^{-4}$ and a batch size of 8.
@@ V Results
This section focuses on the performance of the proposed SQ-GAN evaluated using 500 images (and corresponding SSMs) from the validation set of the Cityscapes dataset [30]. In Section V-A, we illustrate the function and role of the SAMM in selecting features of $\mathbf{x}$ and $\mathbf{s}$ that are semantically relevant. Furthermore, we also show the effect of masking fractions on the final reconstruction. In Section V-B the performance of the proposed SQ-GAN is compared with SOTA compression algorithms like BPG, JPEG2000 and deep-learning based alternatives such as HiFiC [12] and FCC [13]. For the selection of the deep learning-based comparison methods, we considered only those providing the open source code and the model weights. Furthermore, we have not considered methods designed for JSCC, since we consider only source coding.
We also computed the SSM $\tilde{\mathbf{s}}$ of the reconstructed image $\hat{\mathbf{x}}$ and compared it with the SSM $\mathbf{s}$ of the original image, to quantitatively assess the capability of the coding method to preserve some semantics aspects of the image to be transmitted, such as identification of objects like "pedestrians", "traffic signs", etc., the estimation of their shape and of their relative positioning.
@@ V-A SAMM function
Figure 7: Visual representation of the latent tensor selection of the SAMM projected in the image and SSM domain. In both cases the masking has been fixed to $m_{\mathbf{x}}=m_{\mathbf{s}}=0.20$ and the regions considered semantically relevant are shown on the right.
SQ-GAN adopts various techniques to force the model to correctly reconstruct the relevant regions of the image. The weighted loss function based on the semantic classes, the semantic-aware discriminator $D_{disc}^{\mathbf{x}}$, and the data augmentation step described in IV-A, have been designed with this idea in mind. Their scope is to guide SAMM to identify and select the relevant latent vectors $\mathbf{z}_{k}^{\mathbf{x}}$ and $\mathbf{z}_{k}^{\mathbf{s}}$. Since the selection is performed on the latent tensor, to understand the effect of SAMM, we project back from the latent tensor to the image (pixel) domain [21]. Fig. 7 shows such a projection for a specific example.  [^2] The projection shown in figure 7 was implemented by leveraging the convolutional structure of the two encoders that allows to trace back the spatial correlation between the latent tensor and the original frame. The left column displays the original $\mathbf{x}$ and $\mathbf{s}$, whereas the right column indicates the regions associated with the latent vectors deemed most relevant by SAMM. In this example, the masking fractions are chosen as $m_{\mathbf{x}}=m_{\mathbf{s}}=0.20$.
It is immediately evident that $G_{\mathbf{x}}$ and $G_{\mathbf{s}}$ consider different regions as relevant. The SAMM in $G_{\mathbf{s}}$ focuses on regions with the most change in semantic classes. Streets, buildings, and the sky require very few associated latent vectors for their representation. In contrast, the regions of the SSM containing semantically relevant classes are strongly prioritized.
Similarly, the SAMM in $G_{\mathbf{x}}$ shows a strong preference for relevant classes like cars and people. However, it also focuses on areas previously ignored, such as the street and the buildings. Indeed, thanks to the conditioning on the SSM, the sub-network $G_{\mathbf{x}}$ is aware of the location and shape of each relevant object, allowing it to focus on other aspects such as colors and textures. As a result, elements like the streets and the sky still require some latent vectors to accurately reconstruct their colors. However, most of the latent vectors are selected from the truly relevant regions.
@@
[…]
The column on the left shows the original $\mathbf{x}$ and $\mathbf{s}$. The other columns show on top the reconstructed $\hat{\mathbf{x}}$ and the generated SSM obtained from $\hat{\mathbf{x}}$ via the INTERN-2.5 SSModel [22]. The bottom row shows the reconstructed $\hat{\mathbf{s}}$. These examples are obtained by fixing the compression level to a total amount of $0.05$ BPP and by letting the masking fractions $m_{\mathbf{x}}$ and $m_{\mathbf{s}}$ to vary, while keeping their sum constant. In this example, $m_{\mathbf{x}}+m_{\mathbf{s}}=1.1$.
Figure 9: SSM retention evaluated as the mIoU between the true $\mathbf{s}$ and the reconstructed $\hat{\mathbf{s}}$ as a function of the masking fraction $m_{\mathbf{s}}$. As the masking fraction $m_{\mathbf{s}}$ increases the network $G_{\mathbf{s}}$ is able to better reconstruct the SSM. However, the increase of performance reaches a plateau from $m_{\mathbf{s}}\geq 0.2$ ($BPP_{\mathbf{s}}=0.011$BPP).
Figure 10: LPIPS evaluated between $\mathbf{x}$ and $\hat{\mathbf{x}}$ as the masking fractions $m_{\mathbf{s}}$ and $m_{\mathbf{x}}$ vary. By properly selecting the masking fractions it is possible to reduce the redundancy, while preserving a task-specific feature, represented by the LPIPS value
Conversely, the influence of $m_{\mathbf{x}}$ on the overall quality exhibits the expected behavior: increasing $m_{\mathbf{x}}$ enhances image fidelity, particularly for non-relevant details. Relevant features are prioritized and reconstructed effectively even at low values of $m_{\mathbf{x}}$, whereas finer details, such as building windows, are better preserved when $m_{\mathbf{x}}=0.95$. This demonstrates the model’s capability to prioritize semantically relevant features before refining non-relevant aspects of the image.
Original 0.0190.019BPP 0.0380.038BPP 0.0780.078BPP
@@
The impact of $m_{\mathbf{s}}$ on the mIoU is illustrated in Fig. 10. It is useful to notice that, for $m_{\mathbf{s}}\geq 0.55$, the value of mIoU remains almost constant. This means that, above that value, additional latent vectors do not add relevant semantic features. For very low values of $m_{\mathbf{s}}$ the performance is poor, but, as soon as $m_{\mathbf{s}}\geq 0.20$, the model is able to reconstruct $\hat{\mathbf{s}}$ with an acceptable level of semantic retention. To give a term of comparison, the value of $m_{\mathbf{s}}=0.20$ corresponds to a compression of $BPP_{\mathbf{s}}=0.011$ BPP. This means that at this low value of BPP the model is already able to preserve valuable details in the SSM. Increasing $m_{\mathbf{s}}$ beyond the value of 0.55 (equivalent to $BPP_{\mathbf{s}}=0.025$ BPP) does not provide further improvement. We interpret this saturation as follows: beyond this value, all semantically relevant latent vectors have been selected and quantized. Adding other vectors will only increase the amount of redundant information, without improving the reconstructed SSM. This behavior is quite different from what was observed and discussed previously about $m_{\mathbf{x}}$, where even for values close to $1$, the improvements were still visible on the final output.
It is also interesting to see how both $m_{\mathbf{x}}$ and $m_{\mathbf{s}}$ affect $\hat{\mathbf{x}}$. For this purpose, in Fig. 10 we report the value of LPIPS to measure the distance between $\mathbf{x}$ and $\hat{\mathbf{x}}$ and show the importance of semantic masking. As expected, the performance along the $m_{\mathbf{s}}$ axis has a strong influence on the output only for values of $m_{\mathbf{s}}\leq 0.55$. Instead, the influence of $m_{\mathbf{x}}$ is observed along the whole range from 0 to 1. This is consistent with the visual results shown in Fig. 8 and show how, by properly selecting the masking coefficients $m_{x}$ and $m_{s}$, we can strongly reduce the redundancy, while preserving a task-specific feature, represented by the LPIPS value.
@@ V-B Visual Results and Comparisons with SOTA image compression
In this section, the results of the proposed SQ-GAN are compared with the SOTA conventional algorithms BPG and JPEG2000, and with deep-learning-based compression algorithms HiFiC [12] and FCC [13]. A visual comparison between BPG and SQ-GAN is shown in Fig. 11.
The top row represents the reconstructed image $\hat{\mathbf{x}}$ obtained with the proposed SQ-GAN and the associated SSM generated via the INTERN-2.5 SSModel. The bottom row shows the reconstructed image obtained by using the BPG algorithm and the associated generated SSM.
Figure 12: Performance comparison between BPG, JPEG2000 and SQ-GAN in terms of semantic metrics and the classic pixel-by-pixel PSNR.
We notice that SQ-GAN can reach lower values of BPP. In particular, in this case BPG cannot compress images at BPP lower than $0.038$. Furthermore, while BPG uses precious resources to reconstruct semantically irrelevant details such as the windows of the buildings, SQ-GAN focuses on the relevant parts. This is further evidenced by the amount of semantic retention of the generated SSM. For example, the cyclist on the bike is still visible and can be correctly identified by SSModel, at all levels of BPP in Fig. 11. In contrast, the cyclist completely disappears in the generated SSM from the BPG reconstructed image at all levels of BPP in Fig. 11.
@@
[…]
To assess the performance of SQ-GAN, we compare it with established SOTA image encoders, including JPEG2000, BPG, FCC, and HiFiC, using four metrics: FID, LPIPS, Peak Signal-to-Noise Ratio (PSNR), and mIoU. The first three metrics are computed between the original image $\mathbf{x}$ and its reconstruction $\hat{\mathbf{x}}$, while the mean Intersection over Union (mIoU) is evaluated between the original semantic segmentation map $\mathbf{s}$ and the SSM $\tilde{\mathbf{s}}$, generated from $\hat{\mathbf{x}}$ using the SOTA INTERN-2.5 SSModel. The quantitative results are shown in Fig. 12, which plots each metric as a function of BPP, as defined in (4).
We begin by examining how SQ-GAN compares to approaches that are able to compress the image at comparable values of BPP, namely JPEG2000, BPG, and FCC. As evident from Fig. 12, SQ-GAN significantly outperforms these methods on semantic-aware metrics such as FID, LPIPS, and mIoU, which better reflect perceptual and task-related image quality. SQ-GAN is outperformed by BPG and FCC on PSNR, a pixel-level distortion measure, which favors pixel-by-pixel recovery rather than semantic relevance.
It is important to note that SQ-GAN is capable of operating effectively even at very low bitrates. For instance, even at approximately $0.01$ BPP, it preserves meaningful semantic structures, as demonstrated by favorable FID, LPIPS, and mIoU scores. It is also interesting to observe what happens at BPP =0.038=0.038. BPG reconstructs images with a PSNR of 26, outperforming SQ-GAN in terms of pixel fidelity. However, as shown in Fig. 11, BPG fails to preserve important semantic content. In fact, only the first two cars in the scene are detected while the distant pedestrian and the traffic signs are lost. In contrast, SQ-GAN maintains better object-level consistency.
The comparison with the FCC model further highlights the efficiency of SQ-GAN in semantic preservation. For example, FCC achieves an FID=93 at a BPP of $0.061$, whereas SQ-GAN attains the same FID at a BPP of $0.017$, indicating a bitrate reduction factor of approximately $3.6$ times to achieve comparable reconstructed image quality. Similar trends are observed in LPIPS and mIoU, reinforcing the semantic coding advantage of our model.
Finally, we considered the HiFiC model, a High-Fidelity Generative Image Compression method often adopted as a benchmark. We chose not to include HiFiC in Fig. 12, as its operational BPP range begins at $0.142$, which is substantially higher than the range analyzed for the other models in the figure (from $0.015$ to $0.08$ BPP). Working at BPP =0.142=0.142, HiFiC yields superior performance in terms of PSNR=28.34=28.34, FID=35.47=35.47, and LPIPS=0.223=0.223, as expected, because of the much higher BPP. Nevertheless, it is impressive to notice that, in terms of mIoU, HiFiC reaches only a value of $0.565$, i.e., the same value attained by SQ-GAN at BPP =0.018=0.018, which corresponds to a rate reduction factor of about $7.8$ times. This underlines the remarkable semantic compression capability of SQ-GAN, even when compared with high-performing learned codecs at significantly higher bitrates.
@@ VI Conclusions
This work presents SQ-GAN, a new approach for simultaneous encoding of images and their semantic maps, aimed at preserving the image semantic content, even at very low coding rates. The comparison of SQ-GAN with SOTA encoders has been carried out using a set of perceptual performance indicators, like FID, LPIPS, and mIoU, that somehow measure the semantic distortion between the original image (or semantic map) and the reconstructed image (or semantic map). Using these performance indicators, extensive evaluations on the Cityscapes dataset show that SQ-GAN significantly outperforms JPEG2000, BPG, FCC and HiFiC, in terms of rate-(semantic) distortion trade-off. At the same time, SQ-GAN is clearly outperformed by BPG and FCC in terms of conventional pixel-level indicators, such as the PSNR. In summary, SQ-GAN significantly outperforms state-of-the-art image encoders in semantic-aware image compression tasks, whose primary goal is not perfect pixel-level reconstruction, but rather the preservation of semantic contents that are more important for high-level vision tasks.
We tested our method on the autonomous driving use case, but our approach can be applied to many other scenarios, like medical imaging, remote sensing, space exploration, etc., whenever there is a need to strongly compress the images, while retaining as much as possible their semantic content. Nevertheless, also the autonomous driving use case offers useful motivations for using our coding scheme, like reducing the traffic load necessary to send images to a remote server for training autonomous driving models; or strongly compressing images for liability assessment in case of accidents, while preserving critical semantic details (e.g., traffic lights, road signs, pedestrians). Further investigations are in progress to simplify the coding strategy, optimize the shapes of the image and semantic map tensors and generalize the approach to video coding.